OpenStudy (eric_d):

Express in surd form a)tan (-15) d)cot 75 @ganeshie8

ganeshie8 (ganeshie8):

hint : -15 = 30-45

ganeshie8 (ganeshie8):

familiar with angle sum/difference formulas ?

OpenStudy (eric_d):

Not familiar.. But I need to learn these formulas..

ganeshie8 (ganeshie8):

http://oakroadsystems.com/twt/sumdiff.htm#TanSum

ganeshie8 (ganeshie8):

\[\large \tan (A - B) = \dfrac{\tan A - \tan B}{1 + \tan A \tan B}\]

ganeshie8 (ganeshie8):

\[\large \tan (-15) = \tan (30 - 45) = ?\]

OpenStudy (eric_d):

 like this